INFRARED SPECTRA OF CS_{2} TRIMER: OBSERVATION OF AN ISOMER WITH D_{3} SYMMETRY

M. REZAEI, J. NOROOZ OLIAEE, \underline{N. MOAZZEN-AHMADI}, Department of Physics and Astronomy, University of Calgary, Calgary, AB T2N 1N4, Canada; A.R.W. MCKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada.

Infrared spectra of carbon disulfide trimer formed in a pulsed supersonic slit-jet expansion are obtained via direct absorption of a tuneable diode laser in the region of the CS_{2} ν_{3} fundamental band (∼ 1535 cm^{−1}). Two bands sharing the same lower state were assigned to (C^{32}S_{2})_{3}. These correspond to the two infrared active vibrations (a parallel and a perpendicular band) of the constituent CS_{2} asymmetric stretches. The weaker perpendicular band is centered at 1524.613 cm^{−1}. It is shifted by −10.74 cm^{−1} with respect to the free CS_{2} monomer. The parallel band which is significantly stronger is centered at 1545.669 cm^{−1}. It has a vibrational shift of +10.31 cm^{−1}. The transitions with K ≠ 3n and those with J = odd and K = 0 in the ground state are absent in both bands. This rigorously establishes that this isomer of carbon disulfide trimer has D_{3} symmetry. Thanks to the symmetry, only two parameters, the C-C distance and the angle between a monomer axis and the plane containing the C atoms, are required to fully define the structure of the trimer. These were determined to be 3.811 Å and 61.8^{∘}, respectively. In addition, a parallel band corresponding to a single ^{34}S substitution was observed around 1544.46 cm^{−1}. Results from this band are in complete agreement with those of the two bands for (C^{32}S_{2})_{3}.